TOPOLOGICAL VECTOR SPACES OVER VALUED FIELDS II by V.Benekas and A.K.Katsaras, Ioannina, Greece

Abstract. The study of topological vector spaces over valued fields, initiated in [2], is continued. Barrelled spaces, Baire-like spaces, Suprabarrelled spaces and Bornological spaces are investigated. Mathematics subject classification: 46A Keywords: valued fields, barrelled space, Baire-like space, local convergence, bornological space.